OpenStudy (anonymous):

In my lesson it says that 100[1(0)- 0(0)] equals 100(1). But, how can that be if anything times 0 equals 0?

OpenStudy (nikolas):

Can you give more context about your lesson? Assuming this is the same as 100*((1*0)-(0*0)), you are correct, it works out to 100*(0-0)=100*0=0

OpenStudy (anonymous):

The lesson is called determinants of matrices. This is the whole problem. 1[0(2)-2(0)]-10[1(2)-0(0)] + 100[1(0)-0(0)]

OpenStudy (anonymous):

@Nikolas

OpenStudy (nikolas):

I think it is still a mistake - it should be 0 - 20 + 0, giving a final answer of -20. This is what I got when I tried using an online matrix calculator at http://www.bluebit.gr/matrix-calculator/ too.

OpenStudy (anonymous):

yeah that's what I thought. okay, thank you. :-) @Nikolas

OpenStudy (nikolas):

No problem. The world would be a much better place if there were laws against confusing typos in solution books.

OpenStudy (anonymous):

lol yes. now I'm unsure about the whole lesson. smh...